# A mix of long-duration hydrogen and thermal storage enables large-scale electrified heating in a renewable European energy system

### v1

@@ 4.1.1 Geographical scope and grid assumptions
The model includes a network of 28 countries (EU27 excluding Malta and Cyprus plus Norway, United Kingdom and Switzerland). Transmission network capacities are on a net transfer (NTC) basis and fixed to the TYNDP 2022 Reference Grid [entso-e_tyndp_2022] (Figure SI.8, left panel).
@@ 4.1.4 Demand
The demand side of the power system balance consists of an exogenous and an endogenous part. The exogenous part features a historical base electricity demand time series from Open Power System Data [muehlenpfordt_open_2020] that has been corrected for existing electrified heat demand using data from When2Heat [ruhnau_update_2022]. We add incremental industrial electricity demand resulting from projected process electrification until 2050 [neumann_potential_2023] assuming a flat profile and, for computational reasons, we also assume a flat demand profile to represent electric vehicle demand in 2050 based on the size of today’s car fleet. The implicit assumption is that the short-term variations in mobility demand and the flexibility potential of electric vehicles do not interact substantially with long-duration storage requirements beyond the load they add to the system. The left panel of Figure SI.9 gives the resulting (exogenous) annual electricity demand by country.
@@ 4.1.5 Hydrogen
A country’s hydrogen supply can either be met by domestic production using proton exchange membrane (PEM) electrolysis, imports from other countries within, or imports from countries outside the scope of the model. In the case of the latter, imports can arrive by ship or by pipeline. Such external imports are assumed to be possible only under long-term contracts (LTC) that provide a baseload supply with some hourly flexibility of +/- 10%. We assume that shipped hydrogen comes in as cracked ammonia [spatolisano_ammonia_2023]. Import and interconnection capacities are taken from TYNDP 2024, see Figure SI.8, right panel. We assume LTC import prices of 86.8 €/MWh for shipped hydrogen (ammonia + cracking) and 42-51 €/MWh for pipeline imports [entso-e_tyndp_2025]. To use the network, hydrogen generated by electrolysis needs to be compressed, which creates electricity demand. Longer transport distances requires booster compression, resulting in additional electricity demand.
Hydrogen is stored in either salt caverns underground or more expensive pressurized tanks. For both storage options additional compression is necessary. Cavern storage potentials by country are taken from [caglayan_technical_2020] (see Figure SI.10, right panel). Industrial demand for hydrogen is assumed to have a baseload profile and annual quantities are taken from TYNDP 2024 (see Figure SI.9, middle panel). We assume that all other uses of hydrogen, and especially its derivatives, will be covered by additional imports and do not enter the hydrogen balance. Another source of hydrogen demand stems from hydrogen gas turbines that combust hydrogen to produce electricity. The combination of electrolysis, cavern storage and hydrogen turbines form the long-duration electricity storage (LDES) discussed in this paper.
@@ 4.1.6 Heating
The model includes the share of the heating sector that is assumed to be electrified with heat pumps. We assume the use of air-sourced heat pumps. The efficiency of heat pumps varies with the outside temperature, see Figure SI.11, lower panel. For decentralized heat pumps, we set the heat pump capacity exogenously to meet the peak heat demand. We assume that decentralized are equipped with a small buffer storage with a duration of 1.5 hours [roth_power_2024].
@@ 4.2.2 Weather data
We use the largest complete dataset of weather inputs for energy system models that is openly available at the time of writing [antonini_weather-_2024]. We use data from 1940 to 2018. The dataset includes country-level capacity factors for solar photovoltaics, wind onshore and wind offshore as well as inflow data for run-of-river, reservoirs and open pumped-hydro storage. Importantly, it also includes residential and commercial heat demand data for all European countries. The data is based on the ERA-5 reanalysis data of the Copernicus project [hersback_era5_2023] and we use temperature data from the same dataset to generate coefficient of performance time series using the When2Heat tool [ruhnau_update_2022]. We have corrected heat demand and hydro time series for long-run trends [gotske_designing_2024]. We provide details on our approach in SI.XX. Finally, we scale heat demand by annual data from [rozsai_m_jrc-idees-2021_2024]. The resulting annual demands are shown in Figure SI.9, right panel. We note that total demands and implicit heating thresholds reflect today’s building stock. Large-scale insulation and retrofitting could significantly reduce effective useful heating demand in 2050 [zeyen_mitigating_2021].
For the scenarios with central heating networks, we apply the country-wise shares of district heating of total heat demand projected in an optimistic scenario for 2050 by [fallahnejad_district_2024] (Figure SI.10, left panel). For each hour, we apply these shares to the hourly heat demand and assume that the remainder up to 80% is to be supplied by decentral heat pumps, effectively creating two heat demand time series per country.
@@ 4.3 Scenario design
We assume that 80% of space and water heat demand in residential and commercial buildings (not industry) will be covered with electrified heating options, namely decentralized small-scale heat pumps and large-scale heat pumps within district heating networks. These assumptions conservative in some respects and progressive in others.
The assumed share of 80% is surely at the upper end of range of projected electrification shares. A study by the Joint Research Center (JRC) of the European Commission [tsiropoulos_towards_2020] gives an overview of the final energy consumption in the EU28 buildings sector of different scenarios that reach at least 90% emission reduction by 2050. The share of electricity and ambient ranges around 60% for most scenarios, yet reaches up to 80% for others. The leading German energy transition policy research consortium considers various scenarios for the building sector in Germany, in the majority of which decentralized heat pumps and district heating cover more than 80% of the total heat demand by 2045 [luderer_energiewende_2025].

### v2

@@ 4.1.1 Geographical scope and grid assumptions
The model includes a network of 28 countries (EU27 excluding Malta and Cyprus plus Norway, the United Kingdom and Switzerland). Transmission network capacities are on a net transfer (NTC) basis and fixed to the TYNDP 2022 Reference Grid [entso-e_tyndp_2022] (Figure SI.8, left panel).
@@ 4.1.4 Demand
The demand side of the power system balance consists of an exogenous and an endogenous part. The exogenous part features a historical base electricity demand time series from Open Power System Data [muehlenpfordt_open_2020] that has been corrected for existing electrified heat demand using data from When2Heat [ruhnau_update_2022]. We add incremental industrial electricity demand resulting from projected process electrification until 2050 [neumann_potential_2023], assuming a flat profile and, for computational reasons, we also assume a flat demand profile to represent electric vehicle demand in 2050 based on the size of today’s car fleet. The implicit assumption is that the short-term variations in mobility demand and the flexibility potential of electric vehicles do not interact substantially with long-duration storage requirements beyond the load they add to the system. The left panel of Figure SI.9 gives the resulting (exogenous) annual electricity demand by country.
@@ 4.1.5 Hydrogen
A country’s hydrogen supply can either be met by domestic production using proton exchange membrane (PEM) electrolysis, imports from other countries within, or imports from countries outside the scope of the model. In the case of the latter, imports can arrive by ship or by pipeline. Such external imports are assumed to be possible only under long-term contracts (LTC) that provide a baseload supply with some hourly flexibility of +/- 10%. We assume that shipped hydrogen comes in as cracked ammonia [spatolisano_ammonia_2023]. Import and interconnection capacities are taken from TYNDP 2024, see Figure SI.8, right panel. We assume LTC import prices of 86.8 €/MWh for shipped hydrogen (ammonia + cracking) and 42-51 €/MWh for pipeline imports [entso-e_tyndp_2025]. To use the network, hydrogen generated by electrolysis needs to be compressed, which creates electricity demand. Longer transport distances require booster compression, resulting in additional electricity demand.
Hydrogen is stored in either salt caverns underground or more expensive pressurized tanks. For both storage options, additional compression is necessary. Cavern storage potentials by country are taken from [caglayan_technical_2020] (see Figure SI.10, right panel). Industrial demand for hydrogen is assumed to have a baseload profile, and annual quantities are taken from TYNDP 2024 (see Figure SI.9, middle panel). We assume that all other uses of hydrogen, and especially its derivatives, will be covered by additional imports and do not enter the hydrogen balance. Another source of hydrogen demand stems from hydrogen gas turbines that combust hydrogen to produce electricity. The combination of electrolysis, cavern storage and hydrogen turbines forms the long-duration electricity storage (LDES) discussed in this paper.
@@ 4.1.6 Heating
The model includes the share of the heating sector that is assumed to be electrified with heat pumps. We assume the use of air-sourced heat pumps. The efficiency of heat pumps varies with the outside temperature, see Figure SI.11, lower panel. For decentralized heat pumps, we set the heat pump capacity exogenously to meet the peak heat demand. We assume that decentralized heat pumps are equipped with a small buffer storage with a duration of 1.5 hours [roth_power_2024].
@@ 4.2.2 Weather data
We use the largest complete dataset of weather inputs for energy system models that is openly available at the time of writing [antonini_weather-_2024]. We use data from 1940 to 2018. The dataset includes country-level capacity factors for solar photovoltaics, wind onshore and wind offshore as well as inflow data for run-of-river, reservoirs and open pumped-hydro storage. Importantly, it also includes residential and commercial heat demand data for all European countries. The data is based on the ERA-5 reanalysis data of the Copernicus project [hersback_era5_2023] and we use temperature data from the same dataset to generate the coefficient of performance time series using the When2Heat tool [ruhnau_update_2022]. We have corrected heat demand and hydro time series for long-run trends [gotske_designing_2024]. We provide details on our approach in SI.XX. Finally, we scale heat demand by annual data from [rozsai_m_jrc-idees-2021_2024]. The resulting annual demands are shown in Figure SI.9, right panel. We note that total demands and implicit heating thresholds reflect today’s building stock. Large-scale insulation and retrofitting could significantly reduce effective useful heating demand in 2050 [zeyen_mitigating_2021].
For the scenarios with central heating networks, we apply the country-wise shares of district heating of total heat demand projected in an optimistic scenario for 2050 by [fallahnejad_district_2024] (Figure SI.10, left panel). For each hour, we apply these shares to the hourly heat demand and assume that the remainder up to 80% is to be supplied by decentralized heat pumps, effectively creating two heat demand time series per country.
@@ 4.3 Scenario design
We assume that 80% of space and water heat demand in residential and commercial buildings (not industry) will be covered with electrified heating options, namely decentralized small-scale heat pumps and large-scale heat pumps within district heating networks. These assumptions are conservative in some respects and progressive in others.
The assumed share of 80% is surely at the upper end of the range of projected electrification shares. A study by the Joint Research Center (JRC) of the European Commission [tsiropoulos_towards_2020] gives an overview of the final energy consumption in the EU28 buildings sector of different scenarios that reach at least 90% emission reduction by 2050. The share of electricity and ambient ranges around 60% for most scenarios, yet reaches up to 80% for others. The leading German energy transition policy research consortium considers various scenarios for the building sector in Germany, in the majority of which decentralized heat pumps and district heating cover more than 80% of the total heat demand by 2045 [luderer_energiewende_2025].